### v2

# Quasi-aperiodic grain boundary phases of tilt grain boundaries in refractory metals

Abstract
We report new quasi-aperiodic, ground-state structures and phase transitions in 5[001] tilt grain boundaries (GBs) in body-centered cubic (BCC) refractory metals Nb, Ta, Mo, and W. 5 tilt GBs have been extensively investigated over the past several decades, with their ground-state structure—composed of kite-shaped structural units—previously thought to be well understood. By performing a rigorous GB structure search that optimizes the number of atoms in the boundary core, we predict new quasi-aperiodic “split kite” phases analogous to those previously found in GBs in face-centered cubic metals. Our results suggest that complex aperiodic phases of GBs appear to be a general phenomenon. Density functional theory calculations yield similar GB energies and structures for kite and split kite phases in all BCC metals studied. Phase-contrast image simulations of split kites show better agreement with experimental observations, offering an alternative explanation for previous microscopy results and motivating future atomically resolved imaging of the GB structure.
Introduction: Most engineering materials are composed of many differently oriented crystallites, or grains, separated by interfaces called grain boundaries (GBs). These interfaces have structures and properties distinct from the bulk material and can greatly influence material properties [1]. Among the vast space of possible GBs described by the five macroscopic degrees of freedom, a special subset is the coincidence site lattice boundaries, for which the adjoining crystals form interfaces with well-defined periodic unit cells [2]. (210)[001] and (310)[001] tilt GBs with their small periodic units serve as convenient model systems as representative high-angle, high-energy boundaries for both density functional theory (DFT) calculations and empirical potentials, and have been extensively investigated over the past several decades.
The structure of the boundaries is considered well understood and composed of ordered kite-shaped structural units shown in Figure 1b and 1e. This kite-shaped structure was experimentally observed by atomic-resolution scanning transmission electron microscopy (STEM) only in face-centered cubic (FCC) Cu [3] and body-centered cubic (BCC) Fe [4, 5], and inferred through phase contrast imaging in other elemental metals [6, 7, 8, 9, 10, 11, 12, 13]. It was also independently predicted in simulations with empirical potentials [14, 15], and density functional theory (DFT) calculations [16, 17, 18] that used a traditional modeling approach known as the -surface method [19] that joins two misoriented crystals without optimizing the number of atoms in the boundary core.
More recently, improved simulation methodologies that allowed the number of atoms in the GB to change revealed that most empirical embedded-atom method (EAM) potentials of FCC metals predict a completely different ground-state structure, which was named “split kites” [20]. This structure required inserting or removing about half of the plane of atoms in the boundary core. Moreover, split kites exist in multiple energy-degenerate states characterized by different arrangements of their structural units without a well-defined periodic unit cell [21]. Due to its large area, the split kite structure in FCC metals has not been investigated with DFT calculations; however, subsequent studies of [110] symmetric tilt GBs in BCC metals revealed analogous complex aperiodic states that were confirmed by DFT calculations [22] and subsequently validated by direct observations in BCC Fe [23].
Motivated by these studies, we revisit tilt GBs in BCC refractory metals represented by Nb, Ta, Mo, and W. We perform GB structure search using a recently developed Grand canonical Interface Predictor (GRIP) code that optimizes the GB core configuration [24]. We find new aperiodic structures of these tilt boundaries that are analogous to the split kites observed in FCC metals. These complex structures are confirmed using DFT to be the ground states in Ta, Mo, and W, and are at least metastable in the other BCC systems studied.
Results: In this work, we focus on two symmetric tilt (310)[001] and (210)[001] boundaries with misorientation angles around the common tilt axis of and , respectively. The energy of the system is modeled using EAM and modified embedded-atom method (MEAM) interatomic potentials, as well as DFT calculations. Using GRIP, we systematically explore GB structures by uniformly sampling relative translations between the two grains, randomly removing a fraction of atoms from the boundary plane, and dynamically sampling different configurations using empirical potentials [24]. As described in our previous studies [20], we track the GB atomic density, , by calculating
where is the total number of atoms in a region containing the GB (in our case, the entire simulation cell), and is the number of atoms in the perfect crystal plane parallel to the boundary. GRIP also explores larger GB area reconstructions by systematically tiling the smallest periodic cell. For each GB degree of freedom described above, dynamic sampling is performed by running molecular dynamics (MD) simulations at different temperatures to rearrange the atoms and find the optimal GB configuration. During the final stage of the optimization, the structures are relaxed at by a conjugate gradient algorithm to calculate the GB energy as
where and are the total energy and number, respectively, of atoms in the GB region, is the cohesive energy per atom in the bulk structure, and is the area of the GB plane.
Representative results of GRIP searches for two boundaries in Nb and Ta are shown in Figure 1. Each gray circle corresponds to a structure generated during the GRIP search. The GB energy of each structure is plotted versus the fraction of GB atoms . The lowest-energy states at indicated by blue squares correspond to the well-known regular kite GB phases of these boundaries shown in panels (b) and (e), respectively. indicates that regular kite structures can be obtained simply by joining two misoriented half-crystals without inserting or removing atoms. This is, however, not the lowest-energy configuration of either boundary. The ground-state structures found in this work correspond to and for (310) in Nb and (210) in Ta, respectively, which means that 2/7 and 1/6 of the (310) and (210) crystal planes, respectively, have to be inserted to transform the regular kite structures into the ground states. The ground-state structures for Nb (310) and Ta (210) are shown in panels (c) and (f), respectively. The front view of the boundary structure still has the characteristic shape of kites, but in the atomic column corresponding to the tip of the kite, the atoms have different positions resulting in splitting. Following the naming convention established by previous studies in FCC metals that found similar structural features [20, 21], we call the new ground states “split kites.” The side views shown in panels (c) and (f) reveal complex arrangement of the atoms in the BCC split kite structures. In the boundary core, they are not aligned with the (002) atomic planes like in the regular kite structure, but occupy positions in between. The complex core structures in all split kite phases require larger reconstructions and insertion of additional atoms to be properly optimized.
Nb (310) | Nb (210) | Ta (210) | Mo (210) | W (210) | Fe (310) | Fe (210) | ||||||||
kite | split kite | kite | split kite | kite | split kite | kite | split kite | kite | split kite | kite | split kite | kite | split kite | |
EP | 0.888 | 0.870 | 0.893 | 0.881 | 1.357 | 1.331 | 1.998 | 2.100 | 2.427 | 2.593 | 1.393 | 1.532 | 1.414 | 1.473 |
DFT | 1.208 | 1.354 | 1.248 | 1.271 | 1.420 | 1.390 | 2.001 | 1.930 | 2.602 | 2.534 | 1.540 | 1.632 | 1.590 | 1.606 |
The generated BCC split kite phases appear to be aperiodic. Using GRIP, we explore larger area reconstructions of these GBs by tiling the smallest unit cell up to 25 times in the direction along the [001] tilt axis. The nearly degenerate lowest-energy structures for different reconstruction sizes of 1 6, 1 9 and 1 25 are shown in Figure 2a–c. While the structures viewed down the tilt axis appear nearly indistinguishable, the different reconstructions have distinct atomic arrangements along the tilt axis. When viewed from the side, the atoms belonging to the tips of the kites shown in yellow form a periodic wave with a period commensurate with the simulation cell size. Figure 2d shows the GB energy of the different split kite structures as a function of repetition along the tilt axis. The energy converges only for very large reconstructions of 10 and larger, although a structure with a similar low energy can be found for smaller reconstructions (e.g., 1 6 for Ta (210)), which we select for subsequent DFT calculations. Even for very large sizes, the energy plot shows oscillations in . A similar convergence behavior is observed in the values of for the minimum-energy structure at each reconstruction size in Figure 2d.
Split kite phases of 5(210) were identified by empirical potentials in all BCC metals we studied. However, for Mo, W, and Fe, their energies predicted by these models are 5– higher than those of the kite ground-state structure. Representative results are shown in Figure 3 for the Mo (210) tilt GB, and similar results for W and Fe are in Supp. Figure S1 and Supp. Figure S2, respectively [25]. Split kites are a metastable phase of this boundary; they are local minima with respect to and remain stable at finite temperatures (homologous temperature of ) in MD simulations with periodic boundary conditions.
To validate the stability of the split kite structures predicted at by GRIP and demonstrate possible GB phase transitions, we perform finite-temperature MD simulations at temperatures ranging from to 0.7. To model GB phase transitions from kites to split kites, the GBs were terminated at open surfaces following the methodology proposed in Ref. [20]. Open surfaces act as sources and sinks of atoms and effectively introduce a grand-canonical environment in the GB core. The simulations reveal that the split kite phases of all metals are dynamically stable. First-order phase transitions are observed with open-surface boundary conditions as illustrated in Figure 1g.
The use of empirical potentials (see Supp. Table S1) enables more thorough structure exploration using GRIP, including the searches of large-area reconstructions that are currently beyond the reach of DFT calculations. Previous studies have demonstrated the accuracy of empirical potentials in predicting new low-energy structures which were subsequently confirmed by DFT and observed in experiments [22, 23]. However, it is also well known that predicted GB structures and energies are sensitive to the choice of empirical potential [26]. We compare the different predictions of EAM and MEAM potentials in Supp. Figure S3 and Figure S4. With this in mind, we calculate the GB energies of each phase obtained from GRIP using DFT by relaxing a smaller block of atoms extracted from the GB region. Additional details on the methodology are provided in the Supplemental Material [25].
The DFT results summarized in Table 1 confirm split kite phases of 5(210) as the ground state in Ta, Mo, and W, and they have nearly indistinguishable energies from the kite phase for Nb and Fe. The GB energies calculated for the (210) regular kites in this study are consistent with previous reports (Supp. Table S2). None of the regular kite or split kite structures change significantly during the structural relaxation with DFT, and the final structures are presented in Supp. Figure S5.
For split kite phases of the 5(310) boundary, the DFT calculations were only attempted for Nb and Fe, which are the only metals where GRIP searches predicted kite and split kites to have close energies. DFT-predicted split kite GB energies are found to be 6– higher than those of kites. We did not explore other metals because the cells would contain over 700 atoms and the calculations would be too computationally demanding.
Discussion: In this work, we revisit structures of [001] symmetric tilt GBs in refractory BCC metals and study their phase behavior using the recently developed GRIP tool [24]. Contrary to established and well-known structures, we find new ground states and demonstrate multiple GB phases by optimizing the number of atoms in the boundary core and exploring large-area reconstructions. The new BCC split kite phases in [001] tilt GBs can be realized in unit cells of different sizes and appear to be aperiodic (Figure 2), similar to previously reported split kite structures in [001] tilt GBs in FCC metals [20] and other [110] tilt GBs in BCC metals [22]. Our findings provide further evidence that complex aperiodic phases of GBs could be a general phenomenon.
The BCC metals investigated here have different relative stability between the two GB phases and different ground-state structures, highlighting the potential importance of electronic effects. For the reconstructions accessible to DFT, we find that regular kites and split kites have close GB energies, and it is possible that larger area reconstructions could further lower the DFT energies. The existence of metastable split kite phases suggests that under some conditions they may be stabilized by temperature, pressure, or chemical segregation to become the ground states [27].
Structures of GBs in several refractory metals were previously systematically investigated by high-resolution TEM (HRTEM) techniques [8, 9, 11, 10, 12, 13]. Unfortunately, in these earlier studies the GB structure was not observed directly. Instead, it was inferred by matching the simulated intensity patterns derived from structures predicted by atomistic modeling with the observed phase contrast in the HRTEM images. To compare our predicted structures with these previous experimental results, we generate the simulated images for split kite and regular kite structures using the multislice method implemented in abTEM [28]. The simulated images of the two different structures can be clearly distinguished in Figure 4. The wave pattern at the tip of the split kite structure introduces a characteristic broadening in the intensity that is notably absent in the regular kite. The same feature can be seen in the experimental image, suggesting that the split kite structure for Nb (310) may be a better match with previous HRTEM results and that our predicted structures may have already been experimentally observed.
Previous experimental studies have also reported rigid body translations of (002) across the GB plane along the [001] tilt axis [29]. We report similar displacements for kites in Nb, Ta, and Mo in Supp. Table S3, consistent with some [16, 17] but not all [18] ab initio studies. Notably, the lack of translations in our simulated Nb (310) split kite structure, in agreement with experimental observations [9], further supports the possibility that this GB phase was previously observed. Given the sensitivity of the simulated structures and images to the simulation parameters, our results motivate further experimental studies using modern aberration-corrected electron microscopes that can directly resolve atomic positions.
Acknowledgements.
This work was performed under the auspices of the U.S. Department of Energy (DOE) by the Lawrence Livermore National Laboratory (LLNL) under Contract No. DE-AC52-07NA27344. The authors thank Pedro Borges, Colin Ophus, and Omar Hussein for helpful discussions. Computing support for this work comes from the LLNL Institutional Computing Grand Challenge program. An award of computer time was also provided by the INCITE program. This research used resources of both the Argonne and Oak Ridge Leadership Computing Facilities, which are DOE Office of Science User Facilities supported under contracts DE-AC02-06CH11357 and DE-AC05-00OR22725, respectively. T.F. acknowledges support from the U.S. DOE, Office of Science under an Office of Fusion Energy Sciences Early Career Award. All figures are produced using matplotlib [30] and OVITO [31].Author contributions: T.F. conceptualized and supervised the project. E.C. performed the simulations. Both authors analyzed the results and wrote the manuscript. The authors declare no competing financial or non-financial interests.
Data availability: The GRand canonical Interface Predictor (GRIP) tool that generated the data for this work can be found at https://github.com/enze-chen/grip. The data that support the findings of this article are openly available [32].
Supplementary Information for
Quasi-aperiodic grain boundary phases of tilt grain boundaries in refractory metals
Enze Chen1,∗ and Timofey Frolov2,†
1 Department of Materials Science and Engineering, Stanford University, Stanford, CA 94305, USA
2 Materials Science Division, Lawrence Livermore National Laboratory, Livermore, CA 94550, USA
Email:
∗enze@stanford.edu;
†frolov2@llnl.gov
Nb | Ta | Mo | W | Fe | |
EAM | Fellinger et al. [36] | Li et al. [37] | Zhou et al. [38] | Marinica et al. [39] | Proville et al. [40] |
MEAM | Yang and Qi [41] | Park et al. [42] | Hiremath et al. [43] | ||
ADP | Purja Pun et al. [44] | Starikov et al. [45] |
Nb (310) | Nb (210) | Ta (210) | Mo (210) | W (210) | Fe (310) | Fe (210) | ||||||||
kite | ref. | kite | ref. | kite | ref. | kite | ref. | kite | ref. | kite | ref. | kite | ref. | |
EP | 0.888 | 1.117 [46] | 0.893 | 1.257 [46] | 1.357 | 1.351 [47] | 1.998 | 2.032 [48] | 2.427 | 2.500 [49] | 1.393 | 1.442 [45] | 1.414 | 1.451 [45] |
DFT | 1.208 | 1.20 [18] | 1.248 | 1.24 [18] | 1.420 | 1.41 [18] | 2.001 | 2.03 [18] | 2.602 | 2.65 [18] | 1.540 | 1.565 [50] | 1.590 | 1.64 [50] |
SK | 1.354 | — | 1.271 | — | 1.390 | — | 1.930 | — | 2.534 | — | 1.632 | — | 1.606 | — |
Nb (310) | Nb (210) | Ta (310) | Mo (310) | Mo (210) | |
EP relax (shift) | 0.888 | 0.893 | 1.315 | 1.743 | 1.998 |
DFT relax (shift) | 1.208 | 1.248 | 1.457 | 1.705 | 2.001 |
DFT relax (no shift) | 1.188 | 1.262 | 1.453 | 1.810 | 2.066 |
Crystallium (no shift) [18] | 1.20 | 1.24 | 1.44 | 1.73 | 2.03 |
SUPPLEMENTARY METHODS
Grand canonical Interface Predictor (GRIP)
We perform atomic-level optimization of GB structures using the open-source, Python-based GRand canonical Interface Predictor (GRIP) code [24], which rigorously explores structural degrees of freedom (DOF) through dynamic sampling. Bicrystal slabs are automatically generated in the code using the Atomic Simulation Environment (ASE) [51] library, and the orthogonal simulation is oriented such that the -axis is the tilt axis direction, the -axis is the orthogonal in-plane direction, and the -axis is the out-of-plane normal direction. We ensure periodicity in the GB plane (-plane) and at least in the -direction for each slab to minimize cell size effects. There is only one GB in the middle of the simulation cell due to aperiodic boundary conditions in the -direction.
For an individual GB, each iteration of the algorithm has three stages. During the first stage, the initial configuration is created by uniformly sampling a set of parameters specified through the input file. In this study, the algorithm randomly samples a replication of the unit GB cell, randomly translates the upper slab in the -plane, and removes a randomly chosen fraction of atoms from the GB. The simulation box size scales with the replications to maintain in-plane periodicity.
During the second stage, it performs dynamic sampling to optimize the GB structure consistent with the imposed DOF. In this study, we use standard finite-temperature MD simulations using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [52] in the canonical () ensemble with a Langevin thermostat and a time step of as our dynamic sampling technique. The temperature and duration of the MD are also randomly sampled and only the atoms in a region above and below the GB plane (referred to as the “GB region”) are allowed to move freely during the dynamic sampling phase. For this study, we sample a temperature range between 0.2 and 0.7 of the melting temperature of each metal, and a duration up to . Finally, the temperature is quickly ramped down to for . Additionally, there is a chance the algorithm skips the dynamic sampling for one iteration and jumps to the third stage.
In the third stage, each GB structure is fully relaxed at using a conjugate gradient minimization scheme, where atoms in the GB and buffer regions (an addition on each side of the GB region) can move freely. The convergence criteria are or relative energy ( in successive iterations) and for forces, with a maximum of valuations for each criterion. We found this strict criteria necessary for full convergence of the GB structures, particularly when the two grains are displaced along the tilt axis. The algorithm repeats these stages on each processor independently until termination, saving each relaxed structure to disk and periodically deleting duplicates. Duplicates are defined as structures with the same value of and to three decimal places, and the algorithm will keep the structure with a smaller reconstruction and relative translations.
For each relaxed structure, the GB energy, , is computed according to:
as explained in Equation 2 in the Results section of the main manuscript. We also track the fraction of atoms in one plane or GB atomic density, , according to:
as explained in Equation 1 in the Results. Multiple interatomic potential parameterizations are used for each system, as shown in Table S1.
Density functional theory (DFT)
To validate select GB structures, we perform additional density functional theory (DFT) calculations using the Vienna Ab initio Simulation Package (VASP) [53, 54, 55, 56] with projector augmented-wave potentials [57] and the generalized gradient approximation exchange correlation functional of Perdew, Burke, and Ernzerhof [58]. The semi-core states are treated as valence states for Nb, Ta, and Mo (*_pv potential), while Fe and W use the standard potentials. Calculations for Fe are spin-polarized with initial magnetic moments of . We use Monkhorst-Pack [59] -point grids with a density of 7000 points per reciprocal atom (calculated using Pymatgen [60]) and apply Methfessel-Paxton smearing [61] with a width of . The plane wave cutoff energy is and the convergence criteria are set at for energy and for forces.
We create the input structure by extracting a section near the GB region of the optimized structure from GRIP of approximately in thickness (except for Nb (310), which is thick) and adding of vacuum on top. This results in computational cells ranging from 80 atoms (Ta (210) regular kites) up to 662 atoms (Nb (310) split kites). The axes are rescaled to match equilibrium DFT values and atomic positions are fully relaxed while the cell shape and volume are fixed to maintain stresses in the GB plane. The energy of the GB is computed as the difference in total energy of a structure with the GB and a second bulk structure in the same orientation with the same number of atoms and vacuum but without a GB (i.e., a bulk slab), divided by the planar area. This subtraction aims to eliminate the contribution from the two free surfaces in the periodic cell. When the number of atoms between the two structures (GB and bulk) differ, as for split kites, we adjust the energy of the bulk slab by subtracting a DFT-calculated cohesive energy per atom multiplied by the difference in atoms. Mathematically, the GB energy from DFT may be expressed as:
Molecular dynamics (MD)
To study GB phase stability and transitions, we perform finite-temperature MD simulations using methods adapted from previous work [24]. Briefly, we replicate the optimized GB structures in the - and -directions until the simulation cell is around in the -direction and in the -direction along the tilt axis. We freeze the bottom layer of atoms and constrain the top layer to be semi-rigid throughout the simulation, which lasts at least . We use periodic boundary conditions (PBCs) in the -direction and both PBCs and open surfaces with of vacuum in the -direction. We scan a range of temperatures between 0.2 and 0.7 of the melting temperature of each system. For clarity of visualization, we relax all structures at using a conjugate gradient minimization scheme.
Phase-contrast image simulations
To compare our results to previous experimental studies that used high-resolution transmission electron microscopy (HRTEM), we perform phase-contrast image simulations using the multislice algorithm implemented in the abTEM code [28]. The optimized grain boundary structures from GRIP for Nb (310) are used as input structures, replicated as necessary to achieve a lateral distance of at least and a thickness of at least along the tilt axis. We use 64 frozen phonons to approximate atomic vibrations [62] with a standard deviation of atomic displacements of and the infinite potential parametrization from Kirkland [63]. The electron beam energy is set at and the contrast transfer function parameters are (defocus of ), (spherical aberration), and .